Circular No. 6461

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ar IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery,

on Technical Pan films taken with the OCA Schmidt telescope on Aug. 20.05 and 20.08 UT, of a supernova (V ∼ 18.5) located at α = 1^{h}31^{m}52^{s}.40, δ = −2^{o}13′03′′.0 (equinox 2000.0), which is 18′′.9 east and 2′′.1 north of the center of an Sa/Sb face-on galaxy of m_{pg} = 15.3. Nothing appears at the location of SN 1996ar on similar films obtained in 1994 Sept. and Nov., or on the original Palomar Sky Survey exposures. A nearby star (V ∼ 17) has position end figures 50^{s}.00, 11′16′′.5.

M. Turatto and S. Benetti, European Southern Observatory (ESO), report: “We have observed SN 1996ar on Aug. 23.26 UT with the ESO 1.52-m telescope (+ Boller & Chivens spectrograph) at La Silla. The noisy spectrum (range 348.6–1100 nm, resolution 1 nm) classifies SN 1996ar as a type-Ia supernova, ∼ 1 month past maximum light. Broad emissions are measured at 420.0, 489.3, 557.6, 579.1, 623.6, 677.2, 685.7, 782.6, and ∼ 932.8 nm. The expansion velocity deduced from the broad minimum of the Si II 635.5-nm line is ∼ 7800 km/s (the redshift — determined from narrow emission lines of Hα, [N II] 658.4-nm, and [S II] of the parent galaxy spectrum — being 0.06).”

COMET C/1996 B2 (HYAKUTAKE)

Continuation to the ephemeris on IAUC 6412:

1996 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Aug. 25 12^{h}13.^{m}93 −68^{o}01.^{′}2 2.333 2.402 81.^{o}5 24.^{o}6 11.1
Sept. 4 12 53.28 −66 54.4 2.576 2.549 77.2 22.7 11.6
14 13 25.94 −65 59.4 2.818 2.692 72.5 20.9 12.1
24 13 54.19 −65 18.2 3.056 2.832 67.7 19.1 12.4
Oct. 4 14 19.45 −64 50.0 3.287 2.969 63.0 17.5 12.8
14 14 42.65 −64 33.6 3.508 3.103 58.3 15.9 13.1
24 15 04.37 −64 27.7 3.717 3.235 54.0 14.4 13.4
Nov. 3 15 24.94 −64 30.8 3.912 3.364 50.2 13.1 13.7
13 15 44.63 −64 41.9 4.090 3.491 47.0 12.0 14.0
23 16 03.56 −65 00.4 4.251 3.616 44.7 11.1 14.2
Dec. 3 16 21.78 −65 25.5 4.393 3.738 43.4 10.4 14.4
13 16 39.32 −65 57.0 4.516 3.859 43.3 10.1 14.6
23 16 56.13 −66 34.8 4.619 3.979 44.6 10.0 14.8
1996 August 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green